### v1

# Hegel and Modern Topology

Abstract
In this paper we show how some fundamental concepts of modern topology can be understood philosophically in the light of Hegel’s Logic as well how modern topological concepts can provide concrete illustrations of many of the concepts and deductions that Hegel used. This paper can be seen as a continuation of our paper [13] where we argued that the prototypes of many fundamental notions of modern topology were already found in Aristotle’s Physics.
General considerations on Hegel’s Logic
Hegel expresses apparently negative views of mathematics in his published works (specially in the Science of Logic). However we need to balance our judgment with the following extract from his letters (see Wallace’s translation of the Encyclopedia Logic, xiv-xv) which is very revealing:
’I am a schoolmaster who has to teach philosophy, who, possibly for that reason, believes that philosophy like geometry is teachable, and must no less than geometry have a regular structure. But again, a knowledge of the facts in geometry and philosophy is one thing, and the mathematical or philosophical talent which procreates and discovers is another : my province is to discover that scientific form, or to aid in the formation of it’.
’ You know that I have had too much to do not merely with ancient literature, but even with mathematics, latterly with the higher analysis, differential calculus, chemistry, to let myself be taken in by the humbug of Naturphilosophie, philosophising without knowledge of fact and by mere force of imagination, and treating mere fancies, even imbecile fancies, as Ideas.’
Our view is that Hegel is essentially concerned with the phenomenology of consciousness (including forms of social, historical, cultural and religious consciousness) and that the Science of Logic represent an attempt at abstracting the structures and processes found therein. True, there is a sense in which consciousness is left behind at the beginning of the Logic, that it proceeds from where the Phenomenology of Spirit left off, that even pure being is incomprehensible by consciousness. Be this as it may, it is still the case, so we contend, that consciousness is the ultimate source of the Logic. Consider this passage from the Encyclopedia Logic:
In other words, every man, when he thinks and considers his thoughts, will discover by the experience of his consciousness that they possess the character of universality as well as the other aspects of thought to be afterwards enumerated. We assume of course that his powers of attention and abstraction have undergone a previous training, enabling him to observe correctly the evidence of his consciousness and his conceptions.
The following appear to us to be some of the fundamental processes of consciousness which also correspond to certain fundamental mathematical structures or constructions:
  • Synthesis. This is reflected in the gluing of geometric objects, various kinds of covers and specially the sheaf condition.
  • Self-reflection. A system which can represent (partially at least) higher order aspects of itself within itself. A prime example is the original synthetic unity of apperception: I know that I am knowing. This is reflected in recursive and inductive definitions.
  • Return-to-self. This is related to the negation of the negation (double negation). An example is found in temporality: something must pass away to reveal its essence - to ti ên einai, literally ’the what it was to be’.
Let us recall the basic trinary structure of Hegel’s Logic. The Understanding seizes a concept as an immediate whole, but then Dialectical reason remembers the complex process of its genesis and its inward relations and tensions (that it, it becomes a category of objects and morphisms). Speculative reason then replies to Dialectical reason that these relations actually flow in both ways rendering the category again a coherent whole - maybe we can view this as realizing the quotient construction in which we find formal inverses to the relations brought forward by Dialectical reason. Or it is like taking an integral of a certain field density over all possible paths between two points. Thus Speculative reason subsumes into bilateral isomorphisms these directed unilateral relations and again brings peace to the concept, then to be seized again by the Understanding at a higher level.
First look at the Logic of Being
The first section of Hegel’s Logic, the Logic of Being, involves an abstract exploration of the concept of space and the differential geometry used in mathematical physics. Hegel, who taught differential calculus both at high school and university, dedicates a long section of the Logic to the infinitesimal calculus, the notions of which illuminate many other passages of the Logic.
In the section on Being, in the parts involving the One and the Other, Limit, the Infinite, the One, and the successive transitions involved - all this suggests a connection to a universal construction in mathematics: that of a completion. We have a finite structure which is limited, incomplete in some determinate way. For instance a field may not be algebraically closed. Something is missing, something extrinsic that needs to become intrinsic (cf. the process of finite algebraic extensions, adjoining roots, etc.). Yet the finite completion, the finite adjoining of the missing aspects generates in turn its own incompleteness at the next level, and so on. This situation is only overcome by assimilation, by incorporation of this extension process, this outflow, into the structure itself - now invariant under this finite ’passing to the limit’. This ’incorporation’ is itself a kind of higher ’limit’, a limit of limits, so to speak (which has a universal property, or a minimality property). We find similar processes in measure theory (in the construction of the Borel $\sigma$-algebra, regular and complete measures) and analysis (completion and approximation theorems). Hegel states in the Encyclopedia Logic that the determinations of Being are exterior to each other, their process is the passing into another.
The transition for Being-for-self to Quantity recalls the proof of Urysohn’s lemma. The main ingredient is the fact that in a locally compact Hausdorff space $X$ if we have a compact set $K$ contained in an open set $U$ then we can find an open set $V$ such that $K\subset V\subset\overline{V}\subset U$. This expresses $K$’s tendency to overcome its limitation relative to $U$. In the proof of Urysohn’s lemma this process of overcoming the limit is multiplied to countably infinite sets indexed by the rationals and then completed into a continuous function, into continuous quantity. Somehow this relationship to the real numbers is implicit in the abstract concept of a locally compact Hausdorff space. There are analogous situations in which valuations, measures, uniparametric semigroups of automorphisms, emerge. For example for continuous functions with compact support Borel measures emerge from bounded positive linear functionals on spaces of such functions.
Pure Quantity or indifferent Quantity (apeiron) is topological, corresponding to what is implicitly determined by a variable, a sequence, a net (cf. the discussion in [13]). Attraction is aggregation, cohesion, expressed by the open sets of a topology. Repulsion is discreteness . The duality and mutual transition between the continuous and discrete is reflected in the fact that geometric structures give rise to algebraic structures and these in turn have their geometric realizations.
In the initial section of the Greater Logic on Measure there is a suggestion that Measure is a kind of conservation law; that which remains invariant under actual, possible or conceivable change and variation. Measure is related to mathematical physics; Hegel states that a Measure is not pure Quantity but Quantity in relationship to something exterior to itself. Hegel also raises the question of natural vs. artificial units of measure. He introduces gradualness (i.e. continuous change in quantity with surprising jumps in quality) and mentions the Sorites paradox. Hegel’s interpretation might be couched in terms of locally constant sheaves (equivalent to covering spaces): local homogeneity and uniformity is reconciled with global non-triviality. Hegel’s knotted line foreshadowed the modern notion constructible sheaf[11], the ’knots’ corresponding to the stratification (cf. singular points of a variety, bifurcation set, etc.).
The section on Measure concerns also how systems combine into larger systems or split again, etc. Hegel uses chemistry and thermodynamics. It suggests the idea of program or function which can be integrated with different inputs producing different behaviors yet still maintaining its essence.
On Quantity
The indifference of Quality expresses an infinitesimal local Quality around a point. The idea of an open set in which the actual size is completely irrelevant, the local homogeneous quality. This can be understood in terms of a basis of a topology $\tau$: a family of open sets such that each open set $U$ containing a point $x$ contains an element of the basis (containing $x$). The ordinary topology of $\mathbb{R}^{n}$ has a basis given by open balls of radius $\frac{1}{n}$. The Hegelian ’indifference’ is expressed by the fact that we can remove any finite (or even some infinite sequences) of such balls from the basis without affecting the topology defined.
The definition of topology is inseparable from the concept of continuity. And yet the ordinary topology has an element of discreteness, or rather, can be generated by a discrete sequence.
The concept of continuity, in our view, is tied to that of sheaf, in particular to the concepts of gluing or synthesis. In its most rudimentary form it is the condition that any open set must be able to be written as a union $\bigcup_{i\in I}B_{i}$ of open sets of a basis.
Consider a category $C$ of open sets. And let $D:I\rightarrow C$ be a diagram, for instance, consisting of open balls containing a given point. In this case the point is a ’limit’ of the diagram in some sense, but not in the category $C$. But we can define the notion of $c$-limit - the idea here is that the limit point is defined implicitly through the complement of this maximal open set.
Given a diagram $D:I\rightarrow C$ for $I$ a directed set, we say that an open set $V$ in $C$ is a $c$-limit of $D$ if i) for any object $i$ in $I$ we have that $F(j)\cap V\neq 0$ for some $j\prec i$ and any open set $V^{\prime}$ satisfying this condition is contained in $V$. For example, for a sequence of balls centered at the origin the complement of the origin is the $c$-limit.
Hegel’s treatment of repulsion and attraction recalls the separation axioms in general topology. And the concept of connectivity is also directly involved with separation. The extremes are that of the coarse and discrete topologies. In the coarse topology everything is merged together, we cannot separate things. In the discrete topology not only can we separate everything but each point is its own separation (this atomization is a form of alienation that features in many crucial moments of the Phenomenology of Spirit).
Hegel’s ’Units’ are like the proximate species of a genus. In Aristotle genus is a continuum, like matter. The species carve up the genus into pieces (which themselves become further genera). But all species form a spectrum, a continuum in sharing the quality of the proximate genus. All open sets (at least basic ones) should share the same quality. The topology is the quality. In the sheafication construction of a presheaf, a section of the constructed sheaf is essentially a unity, a gluing of local sections, a kind of averaging out into a locally homogeneous quality.
Hegel’s treatment of number in the section of Quantum is noteworthy. Given a non-empty finite set $S$ for every $s\in S$ we have $S=S^{\prime}\cup\{s\}$ with $S^{\prime}\subset S$. This is what Hegel seems to call $s$ being a limit of $S$. It matters not which $s$ is used. Hegel’s whole approach suggests an intuitive continuum-based phenomenological view of number, numbers as connected objects which expresses the internal homogeneity and indifference between objects.
Different total orderings of a finite set are equivalent. For each we have $S=\{n\in N:n\leq max\,S\}$. Orderings are external, imposed. Recall also the external relation of Frege’s equinumerous relation. A number is the set of all sets equinumerous to a given set (the extension of the number) and again the precise choice of this set is irrelevant.
Hegel was certainly prescient about the relationship between geometry and arithmetic, the discrete and the continuous. Modern mathematics has confirmed this. Numbers do not add themselves, they are added from outside. Arithmetical operations are imposed from without, much like the category theoretic view of algebraic structures (for instance a natural number object). We could also wonder if intensive quantity corresponds to viewing the real numbers as a set with an order. Hegel seems to have anticipated in this discussion the axiom of induction in Peano Arithmetic.
Homotopy (in algebraic topology) is the passage of Quantity into Quality. It is a changing of shape and size which preserves - and thus defines - a certain quality.
A simplicial set is an abstraction of a topological space, it is a category theoretic abstraction of a geometric form (i.e. a polytope). But it is a geometric form which contains within itself the process of its own genesis or assemblage (cf. the geometric realisation functor associating a topological space to each simplicial set).
The category of sets $Set$ is an exterior, abstract, discrete concept (Hegel’s concept of number seems to be very set theoretic), but the category of simplicial sets $sSet$ represents a greater cohesion between parts and qualitative structure and determination. The morphisms between objects in a simplicially enriched category instead of being a mere sets become a space.
On Being, Quality and Quantity
Consider a category $C$ with only one object $A$. Is $A$ then an initial or terminal object ? Not necessarily either because $hom(A,A)$ can contain other morphisms besides $id_{A}$ (i.e. we might say it has internal symmetries, it is not absolutely simple). If $A$ is a terminal object then $hom(A,A)=\{id_{A}\}$, but then $A$ is also an initial object. And if $A$ is an initial object then $hom(A,A)$ again must be $\{id_{A}\}$ and so $A$ is also a terminal object. This is one (very simple) illustration of the passage between Being and Nothing and Nothing and Being.
If $C$ is a category with two objects $A$ and $B$ and if $A$ is initial and $B$ is terminal (there is a canonical $t:A\rightarrow B$) then if there is a morphism $f:B\rightarrow A$ then $A$ and $B$ must be isomorphic. This is a simple prototype of the quotient construction. But let $C$ have three objects, initial 00, terminal $1$ and $A$. Then it follows from these hypotheses that the canonical $0\rightarrow 1$ must factor a unique way as $0\rightarrow A\rightarrow 1$. Now $A$ is allowed to have internal symmetries ($hom(A,A)$ need not be $\{id_{A}\}$). And postulating a $1\rightarrow A$ means that ”$A$ has an element”. Thus $A$ is now a Something.
On a very simple level in a category we find already the embodiments of the basic Hegelian concepts: Being-in-self, Being-for-other and Being-for-self. Objects, considered in themselves, without arrows, without relations, are very poor and empty. The only morphism an object $A$ must have is $id_{A}$. We could almost say that $id_{A}$ represents $A$ (in some formulations of a category as a partial monoid we can even identify objects with these morphisms). $id_{A}$ represents the bare minimal immediacy of a quality (or being at this stage). This is being-in-self. Now the existence of a morphism $f:B\rightarrow A$ represents a relation between $A$ and $B$, it represents $A$’s being-for-other, $A$’s being relative to $B$. Likewise $g:A\rightarrow B$ represents $B$’s being relative to $A$. The composition of these two $f\circ g:A\rightarrow A$ is in general non-trivial (distinct from $id_{A}$) abd represents being-for-self, a more complex self-relation which is mediated by relations to the Other.
Much topological intuition and anticipation of modern geometry and logic seems to be found in Hegel’s Science of Logic. For instance Hegel writes: the boundary of a thing both belongs and does not belong to it (we discuss the relationship to the Heyting algebra of open topological sets [13] and the logical significance of the boundary in a topological sense). Points are seen both as limits of lines and generators (in the algebraic and indeed Lie group sense). The pure something in Determinate Being (when we abstract it from its Limit) is likened to a local homogenous quality (open set). The limit is seen as a interface with the environment. Constitution (for instance rational thought is called the constitution of man) implies a barrier to be overcome and already a certain rudimentary invariance under change. Hegel even says points, lines, surfaces have a ”dialectical tension”.
Infinity is the negation of the negation (the connection to the double negation topology in topos theory[12] is worth exploring - it is connected to density and indeed probably to self-similarity and scaling). But there is a simple illustration that might be of some interest. Consider the Heyting algebra of open sets of a topological space (which is in general not a Boolean Algebra so we do not have $A\vee\neg A$ or equivalently $\neg\neg A\rightarrow A)$. The negation of an open set $A$, $\neg A$, is defined as the interior of its complement ($int\,A^{c}$). We have $A\subset\neg\neg A$. But not in general an equality. $\neg\neg A$ here represents the Hegelian infinite, which is essential dynamic, it is about evolving, overcoming barriers. Now look at the open set $(0,1)\cup(1,2)$. If we think of this open interval as having a living, flowing nature, what would the natural tendency be of this open set ? It would be to fill in the whole (the gap $\{1\}$), to cohere and merge into $(0,2)$. But if we calculate $\neg\neg((0,1)\cup(1,2))$ we get precisely $(0,2)$. The open sets which are invariant for the operation of double negation are the regular open sets.
There is more that could be said about the One in terms of infinitesimals, stalks of sheaves and germs of functions and their unfoldings. In particular about how the One expels and generates other Ones. Hegel indicates that he is thinking of Leibniz’s monads, of infinitesimal points.
On an elementary level we can think of the One as a fundamental system of neighbourhoods around a point or equivalently as the germ of a point $p$ which is an equivalence class in a quotient construction. Thus it must have representatives actual open sets $U$ containing $p$. But then $U$ itself generates (or represents) germs for all the points $q$ in $U$. Thus each germ will generate a multiplicity of others. A slightly more advanced way of looking at this is in terms of germs of functions (around a singularity) Each germ of a function determines a universal unfolding which gives rised to qualitative local discontinuities around the point (a catastrophe set). Thus the homogeneous strata = attractions and the germs of submanifolds representing the catastrophe sets represent repulsion. In catastrophe theory one speaks of a ”conflict” of attractors.
The idea of a germ of a point $p$ is quite a good illustration of Being-for-self (the One). The germ of a point is defined as an equivalence class (hence via a quotient construction) of all open neighbourhoods $U$ containing $p$. Thus $U$ and $V$ determine the same germ if $p$ is in $U$ and $V$ and $V$ is contained in $U$. For $p$ to be a germ it needs to enter into relationship with the ambient space, it needs the topology of the space. It cannot be a germ merely in itself. Thus to define the germ we need some open set U (the germ is for-another). This $U$ is posited. But at the same time we can deny and discard this $U$ and replace it by a different one $V$ as long as it is equivalent to $U$ (for instance $V$ is contained in $U$ and contains $p$). This is the moment of negation. That is, the germ is not this, not that, etc. It has abrogated complete determination via any specific open set. The whole net or local basis around $p$ represents the ’bad infinity’. The quotient construction that defines the germs represents finally the return-to-self and a unity.
This germ seen as a One has a ”dialectical tension” which we can associate topologically with the idea of instability. It has to unfold and stabilize in the ambient space but by doing so it reveals its internal contradictions and movement of attraction and repulsion. Maybe there is a connection to symmetry breaking in theoretical physics.
Besides the quotient construction another fundamental construction is that of completion (but completions and limits in general are often defined using quotient constructions). Take the category of open sets of a topological space. Then we can define a category theoretic diagram $U_{1}\rightarrow U_{2}\rightarrow U_{3}\rightarrow....\rightarrow...$ (equivalent to a net, or local basis) which represents a germ at a given point. But, as we already remarked, the limit in the category-theoretic sense does not in general exist. Thus we need to complete the topology by adding infinitesimal neighbourhoods. So if $G(p)$ represents the germ-object and $U$ some neighbourhood we now have canonical morphisms $G(p)\rightarrow U\rightarrow G(p)$ which represents return-to-self.
These germs or infinitesimals have both the nature of discreteness (they are point-like in the sense that for a 2decent” topology two distinct infinitesimals are represented by disjoint open neighbourhoods around their respective points) and continuity. Open sets are obtained from infinitesimals by ”integration”, a kind of synthesis or gluing together of infinitesimals (cf. the Leibniz notation for the integral). Thus the category of Quantity is born. This is also related to the sheafication construction discussed above.
Measure from Hegel to Thom
The section on Measure in the Science of Logic is highly suggestive from the point of view of modern topology. We argue that modern differential topology and singularity theory represents a development of key intuitions and concepts found therein.
The relationship between Quality and Quantity in Measure recalls not only that of a physical field spread out through space-time, but the general concept of a sheaf. The base category is often space, or space-time, or some control space, while the target category is phenomenological, qualitative although having in turn a quantitative aspect. The base is extensive, the fiber intensive. The further development in this section of the Logic - and most specially the concept of Nodal Line - appears to us as a prototype of Thom’s Catastrophe Theory expounded, for instance, in his book Mathematical Models of Morphogenesis. Hegel offers an important insight on the essential co-dependence between the qualitative and quantitative which has consequences for applied mathematics. Strictly determined systems have to be studied qualitatively. But qualitative studies are only meaningful and fruitful in the context of a specific determined object. Thus we must employ qualitative concepts to talk about a sculpture (for instance locally the quality of the curvature, the topology of the surface levels) but a merely global qualitative (i.e. topological) study would identify it with a ball. In order words, the topological exists to study the metrical and the metrical has to be studied using the topological. The same goes for differential equations and dynamical systems. As Hegel wrote: the Quantum is Quantity (in the sense of indifferent quantity) endowed with Quality.
Hegel’s discussion on the Sorites besides being related, as we mentioned, to locally constant sheaves, also invokes phase transitions in modern physics. And is not general relativity a partial realization of Hegel’s project of deriving physics entirely from the properties of space and time ? Hegel mentions the influence on the observer on the experiment (in his discussion of thermodynamics) . The transition to Essence suggests that while applied mathematics is legitimate in its own domain and perspective, deeper scientific knowledge has to be found elsewhere - but this latter (whatever it be) must in all cases involve passing through a thorough understanding of the former.
Hegel’s concept of Type suggests the image of a continuous map $F:U\rightarrow\Phi$ for some phenomenological space $\Phi$ where $U$ represents the allowed variations of the independent parameter. Besides the connection to Thom there is also the more general connection to moduli spaces.
The transition to Essence and modern Physics
Generally speaking modern physics has a very Hegelian structure. In Newtonian mechanics we can start with simple postulates about invariance under Galilean transformations and the conservation of angular momentum ’emerges’ as a consequence of the invariance of inertial frames under rotation. Starting from the invariance of the speed of light in special relativity we get that energy and momentum are united in a elegant way as a single 4-vector (which thus ’emerges’) and that special relativity subsumes classical mechanics as a limit case when $c\rightarrow+\infty$ (Aufhebung). There is also a further process of ’abrogation’ in which concepts first considered static and absolute are then revealed to be relative. Being there (Dasein) is only being there through negation of the ’other’. This is exactly what relativity does by making time, length and mass no longer in-themselves but for-the-other (the observer). In classical quantum mechanics, starting from basic postulates about the wave function, observables and the evolution of the system we obtain that by passing to the classical limit we get an interpretation of the operators and the wave function in terms of classical Hamiltonian mechanics (including the wave function expressed in terms of the action $S$). This is another case of ’Aufhebung’. The uncertainty principle (a surprising result) is also case of emergence. The uncertainty principle interacts with the classical concept of angular momentum to give rise to an entirely novel concept / category: that of spin. Quantum mechanics (and specially quantum field theory) often abrogates previous concepts. Even in Bohm theory the concept of an individual particle is no longer tenable once we consider relativistic versions of quantum theory.
The narrative about the transition between a Newtonian absolute frame of reference and the relativity of referential frames is rich in philosophical consequences. Essence is negativity (this can be represented by the functor that collapses a category into the single category), the abrogation and vanishing of being which paradoxically is the retreat of being into itself, being reflected in itself, shining in itself, the realm of relation and mediation and positness. The Newtonian absolute frame of reference corresponds to the naive immediacy of being. Relativity negates this: this frame of reference does not exist, no frame of reference can be an absolute immediate representation of physics. Thus this negation carries over to all frames of reference, for each one of them relativity says: this is not physics. Physics must stand beyond all frames of reference, the depth beneath the surface appearance. As Gödel defended, relativity entails a certain form of phenomenism or idealism. However the essence must manifest, and from the negation of the absolute claims of reference frames follows that the essence must lie in the relation between the reference frames, that is, the movement and relativity between moments of Being, And in physics this web of relations corresponds exactly to representations of Lie groups (or principal bundles). And we have the further duality between the group itself and its representations which correspond to positing, Appearance, manifestation, Existence. The sections about Identity and Difference at the beginning of Essence are all about relations and specially equivalence relations. Physics can be considered as defined by equivalence classes modulo all possible frames of reference.
Essence as Systems Theory and Holology
We should seek an abstract treatment of the category of Essence (Wesen) whereby Being negates itself to enter more deeply into itself and prepare the way for becoming in and for itself in Concept. The categories of Essence correspond to the categories of classical metaphysics, but they also corresponds to Objective Spirit, to nature. In modern terms the proper abstract treatment of the category of Essence (if it is not to lapse back into Being) must involve among other things a general systems theory adequate for theoretical biology. For it is the living organism which gives us light on the categories of existence (as organism and system) and active substance. Living beings manifest, they have appearance and phenomena. They preserve themselves in their own negation. Biology cannot be reduced to physics and is essentially more than physics in the same way that mediated Essence cannot be reduced to immediate Being. Life is process and mediation. The theory of computation on one side is connected to formal logic. But on the other side it is a crude approximation of biocomputation, the computation of nature. What is a suitable mathematical model for biological systems that cannot be reduced to the general scheme of mathematical physics (sheaves, gauge potentials, jet bundles, PDEs, etc.) ? Biology cannot be reduced to physics because it cannot be described by sections of a bundle which evolve according to a closed manifold in a jet bundle, that is, exhibiting the local causality of PDEs or integral equations based on these assumptions. The ’space’ involved may be such that the stalk at any ’point’ (and the subsequent evolution thereof) depends on the information on the entire space. For instance $X$ can be seen as a groupoid and we can think of the groupoid algebra product which for two elements is defined in terms of the entire groupoid. Let physics correspond to a temporal process $\phi:T\rightarrow\Gamma(S)$ where $\Gamma(S)$ are sections over some bundle-like structure over a ’space’ $X$. This structure forms a certain kind of (commutative) algebraic structure. Then biology corresponds to a temporal process $\beta:T\rightarrow B$ in which $B$ is a (non-commutative) algebraic structure which is not isomorphic to sections of a bundle-like object over any ’space’ $X$. However $B$ may have phenomenological ’representations’ or ’projections’ $\pi:B\rightarrow\Gamma(S_{\pi})$ for $S_{\pi}$ a bundle-like structure over some space $X_{\pi}$. Thus living organisms have physical measurable manifestations but cannot be reduced to such.
Perhaps $\infty$-groupoids might play a role. Here we discuss some of the intuitions involved. Consider points in the plane. Internal equality between two (externally unequal) points $a$ and $b$, written $a=b$, is seen as being witnessed by there being a continuous path from $a$ to $b$. In this way reflexivity, symmetry and transitivity have natural interpretations (in fact we obtain a groupoid or more correctly a 1-groupoid). Note that there can be more than one path between two points. For a space $C$ consisting of two disjoint ’chunks’ $A$ and $B$ there is no path from a point of $A$ to a point of $B$. There are two extreme situations. One is the discrete case in which no two externally distinct points can be connected by a path - this is the classical concept of a set. The other extreme is a connected (technically: 1-connected) space in which we can find a single point $p$ such that $p$ can be connected to any other point in the space. Thus the discrete space has the minimum of internal unity and the contractible space the maximum of internal unity. In general we can have a space with a certain number of connected components (’chunks’). If we view internal equality as an equivalence relation (and thus the space as a setoid) then the number of components is the number of equivalence classes.
We now explain how being connected is only the maximum of unity at the lowest level. Given two points $a$ and $b$ there is a space of all paths from $a$ to $b$ denoted by $Path(a,b)$. Only now in such a space ’points’ are paths and the ’paths’ are deformations between paths (homotopies). So given two paths $p_{1},p_{2}$ in $Path(a,b)$ we have that $p_{1}=p_{2}$ is witnessed by there being a ’path’ (i.e. deformation/homotopy) between paths $p_{1}$ and $p_{2}$. Now the same story is repeated at a higher octave. $Path(a,b)$ itself can possess the extremes of being discrete (set-like) or connected or anything between. From elementary algebraic topology we get examples of spaces which are connected but which for two points $a$ and $b$ there are paths $p_{1}$ and $p_{2}$ in $Path(a,b)$ which cannot be deformed into each other (for instance on a torus). In the case where $Path(a,b)$ is connected for every $a$ and $b$ the space is called contractible but we see now why we could also call it 2-connected. There are also closed paths or ’loops’, elements in $P(a,a)$, and there is no a priori reason why these could not be discrete, not mutually deformable into another. Thus we might add something to our characterization of classical ’set’. If we think of a set as a collection of points or atoms, then these should not have interesting internal structure or symmetries and thus we add the condition that $Path(a,a)$ be path-connected for every $a$: that is, all loops are internally the same. We can now continue this process considering deformations of deformations and so forth. If a deformation is a path between paths in $Path(a,b)$, i.e. an element in $Path(p_{1},p_{2})$ where $p_{1},p_{2}:Path(a,b)$, then we can also consider for $d_{1},d_{2}:Path(p_{1},p_{2})$ elements of $Path(d_{1},d_{2})$, that is to say, deformations of deformations. If we can keep iterating the path construction, then we obtain what is called an $\infty$-groupoid (this is only a rough description) which has a specific notion of $n$-connectivity for each level $n$. What kind of $\infty$-groupoid will have the maximum of connectivity or unity, i.e. how can we define a contractible $\infty$-groupoid ? Clearly by requiring at each level that $Path(x,y)$ for any $x$ and $y$ be path-connected, that is, that we have $n$-connectivity at each level $n$.
Now let us see how $\infty$-groupoids might illustrate the metaphysical categories in Essence as well as the philosophy of nature in Hegel. There is a lowest level of inorganic matter which corresponds to classical sets, thus to 0-discreteness. Then living beings have an organic unity which corresponds to 1-connectedness. Each part of a living body is connected to every other one on the lowest level, but can have quite distinct modes of connection. We can also think of this as there being the same life present in every part but functionally specialized. Next we come to mind which is a higher-order organism in that the same mind is present in every part in an essentially similar way (perhaps related to metacognition or ”self-luminosity”), thus we would postulate both 1- and 2-connectedness for instance. The failure of 3-connectivity could reflect the presence of time or sensation. We could also follow ancient metaphysics such as Plotinus and characterize the nous (seen as a unified space of all forms) as being $n$-connected for all $n$, thus a contractible $\infty$-groupoid, the ontological maximum of unity.
Conclusion: the quest for Notion
Perhaps a key to understanding Hegel’s concept of Notion and the concept of Individual, Particular and Universal and their relationship - is to see how these concepts are embodied in the movements of Spirit in the Phenomenology of Spirit. That is, we view Notion are describing the structure and dynamics of a community of consciousnesses united in different ways into a higher consciousness wherein each individual also is fully identified with this universal consciousness. Although in the first part of Notion Hegel uses the terminology of the formal logic of his day, what he is getting at is much deeper. There is here the mirroring of the microcosm and the macrocosm, the human mind itself is like a community of minds. This suggests the theory of fractals as well as Alain Connes’ non-commutative geometry - for these all concern interface between the local and global. There are two extremes: the global contained fully in the local (as in holomorphic functions) or the local completely emptied of any global information: this is the case of Penrose tilings which cannot be distinguished by considering only local information. Hegel in discussing Objective Notion mentions a system in which the whole is implicit in each particular determination: this is the case of holomorphic functions: the germ at each point completely determines the whole function (on a connected component containing that point) by analytic continutation and the whole completely determines the germ at each point. In general this is a case of a sheaf in which the restriction morphisms are isomorphisms. Holomorphic functions can be called holistic because they are uniquely determined by their germs, the whole is in the part. Change a holomorphic function locally and you change it globally. The global information on the function is contained implicitly in any local restriction or germ (which unfolds via analytic continuation). Thus we can say that the ”same” germ is spread out across the whole domain of the function. The case of holomorphic function corresponds to the global being contained in the particular (or local). The global is thus totally local. But there is also the opposite situation in which the local is emptied of any global information. The global is totally non-local. This is the case, as we mentioned above, for the space of Penrose tilings. It is impossible to distinguish between two tilings by looking at local patterns. This could also be brought into connection to non-locality in quantum theory.
The transition between Essence and Notion, involves the culmination into systems of reciprocal causality. It would be interesting to read this in parallel with Connes’ discussion on how many classical constructions yield equivalence relations or groupoids for which classical analysis and geometry are inadequate - in which a conceptual leap is required by developing tools for a non-commutative concept of space (i.e. a more sophisticated generalization of space which upturns the classical relations between the local and the global). The natural topology on such non-commutative spaces defined as the quotient of an equivalence relation is the coarse topology. So the germ of each point is the space itself, in Hegel’s terms, the Individual is the Universal.
Fractals and fractal calculus may well be worth exploring as a formal hermeneutic tool for Hegel’s Logic. In the Encyclopedia Logic Hegel speaks of Objective Notion as undergoing a division into parts each one which contains the totality.